OpenStudy (anonymous):

HELP ASAP PLZ 1. Choose the correct product of (5x − 3)(5x + 3) 2. Choose the correct simplification of (5x − 6)(3x2 − 4x − 3). 3.Choose the correct simplification of (−4g8h5k2)0(hk2)2. 4.Choose the correct simplification of (2a2b6)3. 5.Compare and contrast, in complete sentences, the difference between the products of a2a4 and (a2)4. Provide the expanded form and product of each.

OpenStudy (mathstudent55):

Part 1. follows the pattern: (a + b)(a - b) = a^2 - b^2

OpenStudy (mathstudent55):

Part 2. Multiply each term of the binomial by each term of the trinomial. Then combine like terms.

OpenStudy (anonymous):

PT 1 25x2 + 30x + 9

OpenStudy (mathstudent55):

Part 3. The zero and 2 are exponents, right? Any number to the zero power equals 1 (except 0^0 which is undefined).

OpenStudy (mathstudent55):

For part 1, use the pattern I gave you above orr use FOIL.

OpenStudy (mathstudent55):

Use FOIL: (5x − 3)(5x + 3) = (5x)(5x) + (5x)(3) - (3)(5x) - (3)(3)

OpenStudy (anonymous):

Choose the polynomial written in standard form. xy2 + 4x4y + 10x2 x4y2 + 4x3y + 10x x4y2 + 4x3y5 + 10x2 x6y2 + 4x3y8 + 10x Question 17 (Multipl